OpenStudy (anonymous):

Which line shows the first error in the solution 5(x-)=15 (1) 5x-25=15 5x-25=15 5x=-10 5x/5=-10/5 x=-2 a.line 2 b.line 3 c.line 4 d.line 5

OpenStudy (pawanyadav):

Error... I think question is incomplete.

OpenStudy (anonymous):

what do you mean

OpenStudy (pawanyadav):

There should be 5(x-5)

OpenStudy (anonymous):

you are right my bad

OpenStudy (pawanyadav):

What do you think about answer....

OpenStudy (anonymous):

i really don't know we never done this in class

OpenStudy (pawanyadav):

OK let's start

OpenStudy (anonymous):

ok what do we do first

OpenStudy (pawanyadav):

As by distribution 5*x-5*5=15

OpenStudy (pawanyadav):

So 5x-25=15

OpenStudy (anonymous):

yeah 5x5x5=15

OpenStudy (pawanyadav):

What

OpenStudy (anonymous):

sorry what were you saying

OpenStudy (pawanyadav):

Do you understand that 5x-25=15

OpenStudy (anonymous):

no

OpenStudy (pawanyadav):

What is 5(x-5) Try using distribution...

OpenStudy (anonymous):

i don' get it

OpenStudy (pawanyadav):

5(x-5)=5(x)-5(5) =5x-25 So you have in question 5(x-5)=15 means 5x-25=15 Now you to remove (-25) from LHS what will you do

OpenStudy (pawanyadav):

I think we have to add 25 both side..

OpenStudy (anonymous):

ok you lost can you break it down slower so i can understand

OpenStudy (pawanyadav):

Don't you understand 5(x-5)=5x-25

OpenStudy (anonymous):

no i don't understand i'm sorry

OpenStudy (pawanyadav):

5(x-5) means first multiply 5 with x minus 5 multiply with 5

OpenStudy (pawanyadav):

Do you sense it now

OpenStudy (pawanyadav):

Sorry I think @pooja195 Can understand you more easily.

OpenStudy (anonymous):

i got 20

OpenStudy (pawanyadav):

No,